Furniture Land surveyed 550 consumers and found that 478 were enthusiastic about a new home dècor it plans to show in its store in High Point.

Construct the 99% confidence interval for the population proportion.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = = x / n = 478 / 550 = 0.869

1 - = 1 - 0.869 = 0.131

Z/2 = Z0.005 = 2.576

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 2.576 * (((0.869 * 0.131) /550 )

Margin of error = E = 0.037

A 99% confidence interval for population proportion p is ,

- E < p < + E

0.869 - 0.037 < p < 0.869 + 0.037

0.832 < p < 0.906

The 99% confidence interval for the population proportion p is : (0.832 , 0.906)